# Reconciling Reality through Simulation: A Real-to-Sim-to-Real Approach for Robust Manipulation

### v1

@@ III-D Teacher-Student Distillation with Co-Training on Real-World Data for Sim-to-Real Transfer
Here the first term corresponds to DAgger training in simulation, while the second term co-trains on real-world expert data. This allows the policy to take advantage of small amounts of high-quality real-world data to bridge the perceptual gap between simulation and real-world scenes and improve generalization compared to only using the data from simulation. We empirically demonstrate (Section III-D) that this significantly increases the resulting success rate in the real world. On a practical note, we refer the reader to Appendix IX for additional details on the student-teacher training scheme that enables it to be successful in the proposed problem setting.
@@ IV Experimental Evaluation
To answer these questions, we evaluate RialTo in eight different tasks, shown in Figure 4 and 8. These include 6-DoF grasping and reorientation of free objects (book on a shelf, plate on a rack, mug on a shelf) and 6-DoF grasping and interacting with articulated objects (drawer and cabinet) on a tabletop and opening a toaster, plate on a rack, putting a cup in the trash in more uncontrolled scenes. More details on the tasks such as their sparse reward functions and randomization setups are presented in Appendix VIII. For each task, we consider three different disturbance levels in increasing order of difficulty (see Appendix VIII for more details):
@@
We conduct our experiments on a Franka Panda arm with the default parallel jaw gripper, using 6 DoF Cartesian end effector position control. For perceptual inputs, we obtain 3D point cloud observations from a single calibrated depth camera. More details on the hardware setup can be found in Appendix X. All of the results in the real world are evaluated using the best policy obtained for each method, we report the average across at least 10 rollouts and the bootstrapped standard deviation. Videos of highlights and evaluation runs are available in the website.
Throughout the next sections, we will evaluate RialTo against the following set of baselines and ablations: 1) Imitation learning from 15 and 50 demos (Section IV-A); 2) No co-training on real-world data (Section IV-B); 3) Co-training on demonstrations in simulation (Section IV-B); 4) RialTo from simulation demos (Section IV-C2); 5) Learning from an untargeted set of simulated assets (Section IV-C1); 6) RialTo without distractors (Section V-A); 7) RialTo without demos (Section V-B)
IV-A RialTo Learns Robust Policies via Real-to-Sim-to-Real
@@
In this section, we aim to understand whether RialTo can solve complex tasks, showing robustness to variations in configurations, disturbances, and distractors. We compare our approach of real-to-sim-to-real RL fine-tuning against a policy trained only on real-world demos via standard imitation learning (BC). We report the results of running RialTo’s pipeline starting from 15 demos collected directly in simulation and co-training with 15 real-world demos during the teacher-student distillation. In Section IV-C2 we show a comparison of running RialTo uniquely on real or sim demos.
@@
Figure 1, 10 and the videos in the website qualitatively show how the resulting policies are robust to many kinds of environment perturbations, including moving the robot, moving the manipulated object and target positions, and adding visual distractors that cause occlusion and distribution shift. The policy rollouts also demonstrate error recovery capabilities, correcting the robot’s behavior in closed loop when, e.g., objects are misaligned or a grasp must be reattempted. This highlights that RialTo provides robustness that does not emerge by purely learning from demonstrations.
Only randomization
Distractors
Disturbances
BC (15 demos) 10 $\pm$ 9%
0 $\pm$ 0%
0 $\pm$ 0%
BC (50 demos) 40 $\pm$ 15%
30 $\pm$ 16%
20 $\pm$ 13%
RialTo (15 demos) 90 $\pm$ 9%
70 $\pm$ 14%
60 $\pm$ 16%
@@
We also compare RialTo against behavior cloning with 50 demonstrations to show that the problem is not simply one of slightly more data. Collecting the 50 demonstrations takes a total time of 1 hour and 45 minutes, which is significantly more than the human effort for RialTo for which we collect 15 demos, in 30 minutes, and build the environment in 15 minutes of active time (see Section VI). Although more data improves the performance of direct imitation learning from 10% to 40%, 0% to 30%, and 0% to 20% for the three different levels of robustness, the results in Table I show that RialTo achieves approximately 2.5 times higher success rate than pure BC, despite using less than one third the number of demonstrations and taking less than half of the human supervision’s time.
IV-B Impact of Co-Training with Real-World Data
@@
IV-C Is Real-to-Sim Transfer Necessary?
@@
IV-C1 Real-to-Sim Transfer of Scenes
Instead of reconstructing assets from the target environment, one could train a policy on a diverse set of synthetic assets and hope the model generalizes to the real-world target scene [12, 21]. While this has shown promising results for object-level manipulation, such as in-hand reorientation  [12], it is still an active area of work for scene-level manipulation and rearrangement  [21]. Moreover, such methods require significant effort in creating a dataset of scenes and objects that enables the learned policies to generalize. Acquiring a controller that can act in many scenes is also a more challenging learning problem, requiring longer wall clock time, more compute, and additional engineering effort to train a performant policy on a larger and more diverse training set.
To probe the benefits of RialTo over such a sim-only training pipeline, we compared the performance against a policy trained using only synthetic assets. Using an amount of time effort roughly comparable to what is required from a single user following our real-to-sim approach (see Section VI), we collected a set of 4 drawers from the Objaverse dataset (see Figure 7). Although this is small compared to the growing size of 3D object datasets, we found it non-trivial to transfer articulated objects into simulation-ready USDs and we leave it as future work. Given these manually constructed diverse simulation scenes, we then trained a multi-task policy using RialTo from 20 demonstrations to open the 4 drawers. See Appendix IX-C for the minor modifications to incorporate multi-task policy learning to RialTo.
@@ IV-C1 Real-to-Sim Transfer of Scenes
IV-C2 Real-to-sim transfer of policies
@@
Figure 6 shows the real-world performance of policies trained using RialTo when starting the RL fine-tuning step using real-world demonstrations as explained in III-C1 against using demonstrations provided directly in simulation. We observe that the performance for both cases is very close. These results show that RialTo successfully learns policies with demonstrations from either source of supervision as long as we keep co-training the policies with real-world data in the teacher-student distillation step. Firstly, this indicates that we do not need to collect both demos in sim and real, but we can run RialTo uniquely from the demos in the real world. Furthermore, this flexibility is a strength of our pipeline, as the ease of acquiring different sources of supervision may vary across deployment scenarios – i.e., one could use policies pretrained from large-scale real-world data or obtain data from a simulation-based crowdsourcing platform.
IV-D Scaling RialTo to In-the-Wild Environments
@@
In this section, we scale up RialTo to more uncontrolled and in-the-wild environments. We test RialTo on three different tasks: open the microwave in a kitchen (also shown in Section IV-A), put a cup in the trash, and bring the plate from the sink to the dishrack. We observe that RialTo scales up to these more diverse scenes and continues to perform significantly better than standard imitation learning techniques. In particular, RialTo brings on average a 57% improvement upon standard imitation learning, see Fig 8.
V Further Analysis and Ablations
Pose Randomization
Distractors
RialTo without distractor training 60 $\pm$ 15%
30 $\pm$ 15%
RialTo with distractor training 100 $\pm$ 0%
70 $\pm$ 15%
@@
V-A Training with Distractors
@@
V-B Comparison to RL from Scratch
Open
Book on
Plate on
Mug on
Open
toaster
shelf
rack
shelf
drawer
RL from scratch with 0 demos 62 $\pm$ 2%
0 $\pm$ 0%
2 $\pm$ 0%
0 $\pm$ 0%
0 $\pm$ 0%
RL fine-tuning from 15 real demos 91 $\pm$ 1%
90 $\pm$ 1%
81 $\pm$ 2%
81 $\pm$ 2%
96 $\pm$ 1%
RL fine-tuning from 15 sim demos 96 $\pm$ 1%
89 $\pm$ 1%
82 $\pm$ 2%
82 $\pm$ 2%
95 $\pm$ 1%
@@
V-C RL from Vision
RialTo’s “inverse distillation” procedure to a compact state-space adds some methodological overhead to the system when compared to the possibility of doing RL fine-tuning directly on visual observations. However, as reported in Appendix Fig. 14, on the task of drawer opening, RL from compact states achieves a 96% success rate after 12 hours of wall-clock time, while RL from vision only achieves a 1% success rate after 35 hours. Hence, inverse distilling to state space is necessary because training RL from vision with sparse rewards is prohibitively slow, motivating the methodology outlined in Section III-C1.
VI User study
We analyzed the usability of RialTo’s pipeline for bringing real-world scenes to simulation. We ran a user study over 6 people, User 6 being an expert who used the GUI before and Users 1-5 never did any work on simulators before. Each participant was tasked with creating an articulated scene using the provided GUI. More precisely, their task was to: 1) scan a big scene, 2) cut one object, 3) scan and upload a smaller object, and 4) add a joint to the scene. From Figure 9, we found that the average total time to create a scene was 25 minutes and 12 seconds of which only 14 minutes and 40 seconds were active work. We also observed that the expert user accomplished the task faster than the rest, and twice as fast as the slowest user. This indicates that with practice, our GUI allows users to become faster at generating scenes. We conclude that doing the real-to-sim transfer of the scenes using the proposed GUI seems to be an intuitive process that is neither time nor labor-intensive when compared to collecting many demonstrations in the real world. We provide more details about the study in Appendix XIII.
@@ VII Limitations and Conclusion
While our use of 3D point clouds instead of RGB enables easier sim-to-real transfer, we require accurate depth sensors that can struggle to detect thin, transparent, and reflective objects. Future work may investigate applying RialTo to train policies that operate on RGB images or RGBD, as our framework makes no fundamental assumptions that prevent using different sensor modalities. We are also limited to training policies for tasks that can be easily simulated and for real-world objects that can be turned into digital assets. Currently, this is primarily limited to articulated rigid bodies, but advancements in simulating and representing deformables should allow our approach to be applied to more challenging objects. Even though we show RialTo works on fast controllers, these are still relatively slow to minimize the sim-to-real gap in dynamics, thereafter there is potential to investigate tasks for which faster controllers are needed. In this work, we consider relatively quasistatic problems, where exact identification of physics parameters is not necessary for the constructed simulation. This will become important as more complex environments are encountered. Finally, as we explain in Section XIV, RialTo currently takes around 2 days of wall-clock time end-to-end to train a policy for each task, this time bottleneck makes continual learning infeasible and understanding how to obtain policies faster with minimal human supervision would be valuable. We expect with more efficient techniques for learning with point clouds and better parallelization, this procedure can be sped up significantly.

### v2

@@ II-D Teacher-Student Distillation with Co-Training on Real-World Data for Sim-to-Real Transfer
Here the first term corresponds to DAgger training in simulation, while the second term co-trains on real-world expert data. This allows the policy to take advantage of small amounts of high-quality real-world data to bridge the perceptual gap between simulation and real-world scenes and improve generalization compared to only using the data from simulation. We empirically demonstrate (Section II-D) that this significantly increases the resulting success rate in the real world. On a practical note, we refer the reader to Appendix VIII for additional details on the student-teacher training scheme that enables it to be successful in the proposed problem setting.
@@ III Experimental Evaluation
To answer these questions, we evaluate RialTo in eight different tasks, shown in Figure 3 and 7. These include 6-DoF grasping and reorientation of free objects (book on a shelf, plate on a rack, mug on a shelf) and 6-DoF grasping and interacting with articulated objects (drawer and cabinet) on a tabletop and opening a toaster, plate on a rack, putting a cup in the trash in more uncontrolled scenes. More details on the tasks such as their sparse reward functions and randomization setups are presented in Appendix VII. For each task, we consider three different disturbance levels in increasing order of difficulty (see Appendix VII for more details):
@@
We conduct our experiments on a Franka Panda arm with the default parallel jaw gripper, using 6 DoF Cartesian end effector position control. For perceptual inputs, we obtain 3D point cloud observations from a single calibrated depth camera. More details on the hardware setup can be found in Appendix IX. All of the results in the real world are evaluated using the best policy obtained for each method, we report the average across at least 10 rollouts and the bootstrapped standard deviation. Videos of highlights and evaluation runs are available in the website.
Throughout the next sections, we will evaluate RialTo against the following set of baselines and ablations: 1) Imitation learning from 15 and 50 demos (Section III-A); 2) No co-training on real-world data (Section III-B); 3) Co-training on demonstrations in simulation (Section III-B); 4) RialTo from simulation demos (Section III-C2); 5) Learning from an untargeted set of simulated assets (Section III-C1); 6) RialTo without distractors (Section IV-A); 7) RialTo without demos (Section IV-B)
III-A RialTo Learns Robust Policies via Real-to-Sim-to-Real
@@
In this section, we aim to understand whether RialTo can solve complex tasks, showing robustness to variations in configurations, disturbances, and distractors. We compare our approach of real-to-sim-to-real RL fine-tuning against a policy trained only on real-world demos via standard imitation learning (BC). We report the results of running RialTo’s pipeline starting from 15 demos collected directly in simulation and co-training with 15 real-world demos during the teacher-student distillation. In Section III-C2 we show a comparison of running RialTo uniquely on real or sim demos.
@@
Figure LABEL:fig:teaser, 9 and the videos in the website qualitatively show how the resulting policies are robust to many kinds of environment perturbations, including moving the robot, moving the manipulated object and target positions, and adding visual distractors that cause occlusion and distribution shift. The policy rollouts also demonstrate error recovery capabilities, correcting the robot’s behavior in closed loop when, e.g., objects are misaligned or a grasp must be reattempted. This highlights that RialTo provides robustness that does not emerge by purely learning from demonstrations.
Only randomization Distractors Disturbances
BC (15 demos) 10 $\pm$ 9% 0 $\pm$ 0% 0 $\pm$ 0%
BC (50 demos) 40 $\pm$ 15% 30 $\pm$ 16% 20 $\pm$ 13%
RialTo (15 demos) 90 $\pm$ 9% 70 $\pm$ 14% 60 $\pm$ 16%
@@
We also compare RialTo against behavior cloning with 50 demonstrations to show that the problem is not simply one of slightly more data. Collecting the 50 demonstrations takes a total time of 1 hour and 45 minutes, which is significantly more than the human effort for RialTo for which we collect 15 demos, in 30 minutes, and build the environment in 15 minutes of active time (see Section V). Although more data improves the performance of direct imitation learning from 10% to 40%, 0% to 30%, and 0% to 20% for the three different levels of robustness, the results in Table I show that RialTo achieves approximately 2.5 times higher success rate than pure BC, despite using less than one third the number of demonstrations and taking less than half of the human supervision’s time.
III-B Impact of Co-Training with Real-World Data
@@
III-C Is Real-to-Sim Transfer Necessary?
@@
III-C1 Real-to-Sim Transfer of Scenes
Instead of reconstructing assets from the target environment, one could train a policy on a diverse set of synthetic assets and hope the model generalizes to the real-world target scene [12, 21, 62]. While this has shown promising results for object-level manipulation, such as in-hand reorientation  [12], it is still an active area of work for scene-level manipulation and rearrangement  [21]. Moreover, such methods require significant effort in creating a dataset of scenes and objects that enables the learned policies to generalize. Acquiring a controller that can act in many scenes is also a more challenging learning problem, requiring longer wall clock time, more compute, and additional engineering effort to train a performant policy on a larger and more diverse training set.
To probe the benefits of RialTo over such a sim-only training pipeline, we compared the performance against a policy trained using only synthetic assets. Using an amount of time effort roughly comparable to what is required from a single user following our real-to-sim approach (see Section V), we collected a set of 4 drawers from the Objaverse dataset (see Figure 6). Although this is small compared to the growing size of 3D object datasets, we found it non-trivial to transfer articulated objects into simulation-ready USDs and we leave it as future work. Given these manually constructed diverse simulation scenes, we then trained a multi-task policy using RialTo from 20 demonstrations to open the 4 drawers. See Appendix VIII-C for the minor modifications to incorporate multi-task policy learning to RialTo.
@@ III-C1 Real-to-Sim Transfer of Scenes
III-C2 Real-to-sim transfer of policies
@@
Figure 5 shows the real-world performance of policies trained using RialTo when starting the RL fine-tuning step using real-world demonstrations as explained in II-C1 against using demonstrations provided directly in simulation. We observe that the performance for both cases is very close. These results show that RialTo successfully learns policies with demonstrations from either source of supervision as long as we keep co-training the policies with real-world data in the teacher-student distillation step. Firstly, this indicates that we do not need to collect both demos in sim and real, but we can run RialTo uniquely from the demos in the real world. Furthermore, this flexibility is a strength of our pipeline, as the ease of acquiring different sources of supervision may vary across deployment scenarios – i.e., one could use policies pretrained from large-scale real-world data or obtain data from a simulation-based crowdsourcing platform.
III-D Scaling RialTo to In-the-Wild Environments
@@
In this section, we scale up RialTo to more uncontrolled and in-the-wild environments. We test RialTo on three different tasks: open the microwave in a kitchen (also shown in Section III-A), put a cup in the trash, and bring the plate from the sink to the dishrack. We observe that RialTo scales up to these more diverse scenes and continues to perform significantly better than standard imitation learning techniques. In particular, RialTo brings on average a 57% improvement upon standard imitation learning, see Fig 7.
IV Further Analysis and Ablations
Pose Randomization Distractors
RialTo without distractor training 60 $\pm$ 15% 30 $\pm$ 15%
RialTo with distractor training 100 $\pm$ 0% 70 $\pm$ 15%
@@ IV Further Analysis and Ablations
IV-A Training with Distractors
@@
IV-B Comparison to RL from Scratch
Open Book on Plate on Mug on Open
toaster shelf rack shelf drawer
RL from scratch with 0 demos 62 $\pm$ 2% 0 $\pm$ 0% 2 $\pm$ 0% 0 $\pm$ 0% 0 $\pm$ 0%
RL fine-tuning from 15 real demos 91 $\pm$ 1% 90 $\pm$ 1% 81 $\pm$ 2% 81 $\pm$ 2% 96 $\pm$ 1%
RL fine-tuning from 15 sim demos 96 $\pm$ 1% 89 $\pm$ 1% 82 $\pm$ 2% 82 $\pm$ 2% 95 $\pm$ 1%
@@ IV-B Comparison to RL from Scratch
IV-C RL from Vision
RialTo’s “inverse distillation” procedure to a compact state-space adds some methodological overhead to the system when compared to the possibility of doing RL fine-tuning directly on visual observations. However, as reported in Appendix Fig. 13, on the task of drawer opening, RL from compact states achieves a 96% success rate after 12 hours of wall-clock time, while RL from vision only achieves a 1% success rate after 35 hours. Hence, inverse distilling to state space is necessary because training RL from vision with sparse rewards is prohibitively slow, motivating the methodology outlined in Section II-C1.
V User study
We analyzed the usability of RialTo’s pipeline for bringing real-world scenes to simulation. We ran a user study over 6 people, User 6 being an expert who used the GUI before and Users 1-5 never did any work on simulators before. Each participant was tasked with creating an articulated scene using the provided GUI. More precisely, their task was to: 1) scan a big scene, 2) cut one object, 3) scan and upload a smaller object, and 4) add a joint to the scene. From Figure 8, we found that the average total time to create a scene was 25 minutes and 12 seconds of which only 14 minutes and 40 seconds were active work. We also observed that the expert user accomplished the task faster than the rest, and twice as fast as the slowest user. This indicates that with practice, our GUI allows users to become faster at generating scenes. We conclude that doing the real-to-sim transfer of the scenes using the proposed GUI seems to be an intuitive process that is neither time nor labor-intensive when compared to collecting many demonstrations in the real world. We provide more details about the study in Appendix XII.
@@ VI Limitations and Conclusion
While our use of 3D point clouds instead of RGB enables easier sim-to-real transfer, we require accurate depth sensors that can struggle to detect thin, transparent, and reflective objects. Future work may investigate applying RialTo to train policies that operate on RGB images or RGBD, as our framework makes no fundamental assumptions that prevent using different sensor modalities. We are also limited to training policies for tasks that can be easily simulated and for real-world objects that can be turned into digital assets. Currently, this is primarily limited to articulated rigid bodies, but advancements in simulating and representing deformables should allow our approach to be applied to more challenging objects. Even though we show RialTo works on fast controllers, these are still relatively slow to minimize the sim-to-real gap in dynamics, thereafter there is potential to investigate tasks for which faster controllers are needed. In this work, we consider relatively quasistatic problems, where exact identification of physics parameters is not necessary for the constructed simulation. This will become important as more complex environments are encountered. Finally, as we explain in Section XIII, RialTo currently takes around 2 days of wall-clock time end-to-end to train a policy for each task, this time bottleneck makes continual learning infeasible and understanding how to obtain policies faster with minimal human supervision would be valuable. We expect with more efficient techniques for learning with point clouds and better parallelization, this procedure can be sped up significantly.